OpenStudy (anonymous):

At a local plant sale, Janice was selling Golden Barrel Cactus plants and baby Saguaros. During the sale, she sold 5 times as many Saguaros ($6 each) as Golden Barrels ($10 each). She made a total of $720 during the plant sale. How many Saguaros did Janice sell? A. 18 B. 40 C. 72 D. 90

OpenStudy (anonymous):

you can write an equation to solve this

OpenStudy (anonymous):

i need the answer @mathessentials

OpenStudy (anonymous):

5*s + g = $720

OpenStudy (anonymous):

whats the answer? @mathessentials

OpenStudy (anonymous):

set into s, and g the price of the vegetables :) write equation here

OpenStudy (anonymous):

5*s + g = $720

OpenStudy (anonymous):

i still dotn understand what is the answer? @mathessentials

OpenStudy (anonymous):

we can make two equations is that acceptable?

OpenStudy (anonymous):

wait 5*$6*x + $10*x = $720 how does this work? well, every sougaru costs $6, so a sougaro is $6*x in the eq. (similarily, every goldencactus cost $10, so $10*x from those.) also,i say that the sougaro ($6*x from total income) is 5 times as often OK so that means that whenever there is one golden($10*1) there is ($6*5) or 5 sougarus, right ?? I sell one golden, I get 5 sougarous at the same time to have the total balance as in the problem.

OpenStudy (anonymous):

5*$6*x + $10*x = $720 i can pull the unknown x out: x(5*$6+$10) = $720 x($30+$10) = $720 x($40) = $720 divide by $40 x = $720/$40 = 18

OpenStudy (anonymous):

so she sold 18 golden cactus, and if that is so, then she also sold 5*golden with it ; ) so that would be 18*5 = 90 sougarous.

OpenStudy (anonymous):

we can check if thats correct. 90 sougarous for $6/pc + 18 golden cactus for $10/pc ------------------------- ?? $540 $180 ----- $720, yep that's it!

OpenStudy (anonymous):

@jay45

OpenStudy (anonymous):

the way I solved it was by getting ONE UKNOWN for both vegetables. by doing that, I was able to determine that one unknown. then, I can calculate every separate vegetable.

OpenStudy (anonymous):

what I did was basically to say that, there are just a few possible sellings, if you know one is sold 5 times the other:

OpenStudy (anonymous):

you can change the total revenue, and the amounts will change automatically from 5:1 to 10:2, 15:3 and the price will also be adjusted, since I put $ next to each veg :)